# Minimax and adaptive estimation of general linear functionals under sparsity

### v1

@@ Abstract
We study the problem of estimating a linear functional $\eta^{\intercal}\theta$ of a high-dimensional sparse mean vector $\theta$ with an arbitrary loading vector $\eta$ under symmetric noise with exponentially decaying tails, with Gaussian noise as an important example. We first establish the nonasymptotic minimax rate in the oracle setting with known sparsity level $s$. This rate explicitly depends on the structure of $\eta$, sparsity level $s$, and tail parameter of noise. We then develop an adaptive estimator that does not require knowledge of $s$ and prove its optimality, showing that the cost of adaptation is at most logarithmic in $s$. Our analysis for arbitrary loadings uncovers a new phase transition in minimax estimation that does not arise under homogeneous loadings. In addition, we extend the minimax theory to non-symmetric noise settings and to hypothesis testing, and we further explore the estimation with unknown noise levels.
@@ 1 Introduction
$L(\theta)=\eta^{\intercal}\theta=\sum_{j=1}^{d}\eta_{j}\theta_{j},$ (2)
based on the observations $y_{1},\ldots,y_{d}$, where $\eta=(\eta_{1},\ldots,\eta_{d})\in\mathbb{R}^{d}$ is the constant loading vector. Because coordinates with zero loadings can be dropped, we assume that all $\eta_{j},j=1,\ldots,d$ are nonzero and sorted such that $|\eta_{1}|\geq|\eta_{2}|\geq\cdots\geq|\eta_{d}|>0$.
@@
for some integer $s\in\{1,\ldots,d\}$. Such a sparsity assumption arises in many applications, including spectroscopy, astronomy, and interferometry [12]. In these applications, the observed signal is typically close to zero with a few rare spikes, often described as a nearly black object. Most theoretical work has focused on Gaussian or sub-Gaussian noise with known $\sigma$, although in practice other noise distributions may also be relevant.
In this paper, we denote by $\mathcal{P}_{j}$ the distribution of noise $\xi_{j}$ for $j=1,\ldots,d$. For convenience, denote by $\mathcal{P}_{\xi}=(\mathcal{P}_{1},\ldots,\mathcal{P}_{d})$ the joint collection of noise distributions. We denote by $\mathbb{P}_{\theta,\mathcal{P}_{\xi}}$ the distribution of $(y_{1},\ldots,y_{d})$ when the signal is $\theta$ and the joint distribution of the noise variables is $\mathcal{P}_{\xi}$. We also write $\mathbb{E}_{\theta,\mathcal{P}_{\xi}}$ for the corresponding expectation. The classical Gaussian sequence model corresponds to the case where the noise variables $\xi_{i}$ are standard Gaussian, i.e., $\mathcal{P}_{1}=\cdots=\mathcal{P}_{d}=\mathcal{N}(0,1)$, in which case we write $\mathcal{P}_{\xi}=\mathcal{N}^{\otimes}$.
Beyond the Gaussian setting, we study a broader family of $\mathcal{P}_{\xi}$ that retain symmetry and fast-decaying tails. Specifically, we introduce the following class.
@@
Distributions in $\mathcal{G}_{\alpha,\tau}$ exhibit sub-Weibull tails and have recently attracted considerable attention [11, 24, 21]. The additional symmetry condition is important for our analysis. The class parameter $\alpha$ specifies the tail behavior and parameter $\tau$ specifies the spread of the tail. In particular, when $\alpha=2$, the class reduces to a sub-Gaussian class; when $\alpha=2$ and $\tau=2$, it includes the standard Gaussian distribution.
We denote by $\mathcal{G}_{\alpha,\tau}^{\otimes}$ the class of (independent) product distributions on $\mathbb{R}^{d}$ whose marginals all belong to $\mathcal{G}_{\alpha,\tau}$. In other words, $\mathcal{G}_{\alpha,\tau}^{\otimes}=\{\otimes_{j=1}^{d}P_{j}:P_{j}\in\mathcal{G}_{\alpha,\tau},~~j=1,\ldots,d\}$. We assume the noise distribution $\mathcal{P}_{\xi}$ lies in $\mathcal{G}_{\alpha,\tau}^{\otimes}$ with fixed $\alpha$ and $\tau$ but the marginals of $\mathcal{P}_{\xi}$ may differ and remain unknown.
@@
The minimax rate of estimation under the Gaussian sequence model has been extensively studied in the literature; see, for example, [15, 4, 5, 13, 14, 22]. The most closely related work to ours is [9], who established the nonasymptotic minimax rate for estimating the linear functional with a homogeneous loading vector (i.e., $\eta_{j}=1$ for all $j$) in the Gaussian sequence model. Specifically, for estimating the sum $\sum_{j=1}^{d}\theta_{j}$, they show that
@@
$\sigma^{2}s^{2}\log\!\left(1+\frac{d\log d}{s^{2}}\right),$
which is attained by a Lepski-type estimator. While these theoretical works on the homogeneous loading case yield valuable insights into the roles of sparsity and dimensionality, they do not directly extend to linear functionals with heterogeneous loadings (i.e., unequal $\eta_{j}$’s).
In this paper, we study minimax estimation of a linear functional with an arbitrary loading vector $\eta$ and with any noise distribution $\mathcal{P}_{\xi}\in\mathcal{G}_{\alpha,\tau}^{\otimes}$. Before proceeding, we present two examples where the loadings $\eta_{j}$’s are heterogeneous and the results of [4, 9, 10] do not apply.
@@
Let $\mathcal{H}$ be the class of square integrable functions on $[0,1]$. Suppose $f\in\mathcal{H}$ is unknown and $Y$ is an observation from the white noise model
$dY(t)=f(t)dt+\sigma dW(t)$
where $W(t)$ is a standard Brownian motion. It is often of interest to estimate the integral $\int fg$ for some $g\in\mathcal{H}$. To see the connection to our problem, let $\{\varphi_{i}(t)\}$ be any orthonormal basis for $\mathcal{H}$ and define
$y_{j}=\int_{0}^{1}\varphi_{j}dY,~~\theta_{j}=\int_{0}^{1}f\varphi_{j},~~\xi_{j}=\int_{0}^{1}\varphi_{j}dW,~~\text{ and }\eta_{j}=\int_{0}^{1}\varphi_{j}g.$
It then follows that $\{\xi_{i}\}$ are independent standard normal r.v.s and $\int fg=\sum_{j}\eta_{j}\theta_{j}$. If it is believed that $\theta_{i}=0$ for $i>d$ for some large integer $d$ and $\{\theta_{i}\}_{1}^{d}$ has no more than $s$ nonzero components, this model coincides with (1) and the estimation reduces to estimating $L(\theta)$ in (2). For further discussions, we refer the reader to [18].
@@
Given $n$ covariate vectors $x_{i}\in\mathbb{R}^{d}$, suppose $n$ observations are $z_{i}=x_{i}^{\prime}\beta+\sigma\varepsilon_{i}$ where $\beta\in\mathbb{R}^{d}$ is the unknown parameter and $\varepsilon_{i}$ are i.i.d. samples from $\mathcal{N}(0,1)$. A common goal is to predict the response for a new covariate vector $x_{0}$ by estimating $x_{0}^{\prime}\beta$. To see the connection to our problem, suppose the design matrix $\mathbf{X}=[x_{1}^{\prime},\ldots,x_{n}^{\prime}]^{\prime}$ has rank $d$ and admits the singular value decomposition (SVD) that $\mathbf{X}=\sum_{j\leq d}b_{j}u_{j}v_{j}^{\prime}$ with singular values $b_{j}>0$, left singular vectors $u_{j}\in\mathbb{R}^{n}$, and right singular vectors $v_{j}\in\mathbb{R}^{d}$. For any $j\leq d$, define the transformed variables and parameters as
$y_{j}=\sum_{i\leq n}u_{j,i}z_{i},~~\theta_{j}=b_{j}v_{j}^{\prime}\beta,~~\xi_{j}=\sum_{i\leq n}u_{j,i}\varepsilon_{i},\text{ and }~~\eta_{j}=b_{j}^{-1}v_{j}^{\prime}x_{0}.$
It follows that $y_{j}=\theta_{j}+\sigma\xi_{j}$, $\xi_{j}$ are i.i.d. from $\mathcal{N}(0,1)$, and $x_{0}^{\prime}\beta=\sum_{j=1}^{d}\eta_{j}\theta_{j}=L(\theta)$. If it is assumed that $\beta$ has no more than $s$ nonzero coefficients w.r.t. the basis $\{v_{j}\}$, then $\theta\in\Theta_{s}$. In this case, the model coincides with the sequence model in (1), and the prediction reduces to estimating $L(\theta)$. This setting differs from the vast literature on inference for low-dimensional parameters in high-dimensional models [31, 29, 17], where sparsity is imposed on $\beta$ rather than $\theta$ and the results on minimax optimal inference are restricted to special types of loadings; we refer to [2, 6] for more details.
Contributions. We first investigate the estimation problem in the oracle setting where the sparsity level $s$ is known, and establish the nonasymptotic minimax rate that reflects the heterogeneity in the loading vector. Specifically, we introduce in Equation 5 an important quantity $\lambda_{o}$ that depends on the triplet $(s,\eta,\alpha)$ and we use it to express the minimax rate as $\Phi_{o}(s;\eta)$ in Equation 8. We then study the optimal rate for adaptive estimation with unknown $s$. To obtain sharp minimax rates and adaptive rates, we need to characterize how the structure of the loading vector affects estimation. Unlike prior studies restricted to homogeneous loadings, our theory covers arbitrary loading vectors, which allows us to reveal new phase transition phenomena. The derivation of both upper and lower bounds is nontrivial and requires new technical tools.
Our main contributions are summarized as follows:
  • Upper bounds. The optimal estimators developed in [9, 10] for the homogeneous case are simple (see eq. 4) but do not extend to general loading vectors. With heterogeneous loadings, the difficulty is that the optimal treatment of each term $\eta_{j}\theta_{j}$ in $L(\theta)$ depends on the magnitude of its loading (i.e., $|\eta_{j}|$). To address this, we propose new estimators that adapt to the loadings: coordinates $\theta_{j}$ with large absolute loadings $|\eta_{j}|$ are estimated directly by $y_{j}$, while those with small absolute loadings are handled by thresholding. The threshold is carefully designed to reflect the structure of the vector $\eta$ and the tail behavior of the noise distribution. This differential treatment, together with the choice of the threshold, is crucial for achieving minimax optimality with general $\eta$.
  • Lower bounds. The lower bound arguments for homogeneous loadings in [9, 10] are based on a uniform prior distribution on the set of exact $s$-sparse vectors (i.e., $\|\theta\|_{0}=s$) with equal nonzero coordinates. This construction does not accommodate heterogeneous loadings, since it ignores differences in magnitudes, and therefore fails to provide sharp lower bounds that match the upper bounds. In addition, their analysis is confined to Gaussian noise and does not account for other tail behaviors of the noise distribution (which are governed by the parameter $\alpha$ in $\mathcal{G}_{\alpha,\tau}$). To overcome these limitations, we adapt the random sparsity prior in [8], originally developed for signal detection in Gaussian sequence models, and extend it to construct the least favorable points for linear functional estimation with broader noise families.
  • Adaptiveness and robustness. When the sparsity level $s$ is unknown, we propose an adaptive estimator inspired by Lepski’s method, whose cost of adaptation is at most a logarithmic factor in $s$. We establish its rate as the adaptive rate under a mild condition, and verify this condition in some examples. Moreover, our framework accommodates unknown noise distributions that need not be identical across dimensions, which highlights the robustness of our method compared to traditional approaches that focus solely on i.i.d. Gaussian noise.
@@
We also study the case where the noise is not necessarily symmetric, and show that the lack of symmetry can lead to a larger minimax rate in certain examples. In addition, we consider extensions to the setting with unknown noise variance $\sigma^{2}$ and to hypothesis testing for the linear functional.
@@
Organization: The rest of the paper is organized as follows. In Section 2, we study minimax estimation of the linear functional $L(\theta)$ under symmetric sub-Weibull noise when the sparsity level $s$ is known. Section 3 considers the more realistic case where $s$ is unknown and establishes the optimal adaptive rate for estimating $L(\theta)$. In Section 4, we extend the analysis to general (not necessarily symmetric) sub-Weibull noise with known sparsity. Section 5 provides analytically tractable examples of the derived rates. Extensions and future directions are discussed in Section 6 and Section 7, respectively. The last section contains proofs of the lower bounds.
@@ 2 Minimax estimation with known sparsity
We first verify that $\beta$ is well defined. Indeed, the left-hand side of Equation 5 is continuous and strictly decreasing (see Lemma 11 in the appendix), diverges to $\infty$ as $\beta\to$-$\infty$, and converges to 0 as $\beta\to+\infty$. It follows that Equation 5 admits a unique solution $\beta$ and we have $\beta_{+}=\beta\vee 0\geq 0$. When $\alpha=2$, Equation 5 coincides with Equation (5) in [8], so our formulation extends their definition from Gaussian noise to symmetric sub-Weibull noise. The parameter $\lambda_{o}$ captures the dependence of the linear functional estimation problem on the three key components: the sparsity level $s$, the loading vector $\eta$, and the noise tail parameter $\alpha$. $\lambda_{o}$ plays a crucial role in characterizing the minimax rate, as it appears both in the expression of our lower bound and in the construction of our optimal estimator.
@@
and the estimator $\hat{L}_{s}$ is a minimax rate optimal estimator.
@@
(1). [Lower bound] For $\alpha>0$, there exist some constants $c,\tau_{\alpha}>0$ such that for any $\tau\geq\tau_{\alpha}$ and any integer $s\in[d]$, it holds that
@@
(2). [Upper bound] For the estimator $\hat{L}_{s}$ defined in (7) and $\alpha,\tau>0$, there exists some constant $C>0$ such that
@@
Theorem 1 may appear similar to Theorem 1 in [8], who studied the minimax separation distance in the heteroscedastic Gaussian sequence model for testing $H_{0}:\;\theta=0$ against $H_{1}:\;\|\theta\|_{t}\geq\varepsilon$ where $\|\theta\|_{t}=\left(\sum_{j=1}^{d}|\theta_{j}|^{t}\right)^{1/t}$ for $t\geq 1$. As discussed in their work, the results on testing connect closely to those on minimax estimation of $\|\theta\|_{t}$, but the case $t=1$ does not cover ours. The functional $\|\theta\|_{1}=\sum_{j=1}^{d}|\theta_{j}|$ is nonlinear and sign-invariant, whereas our target $L(\theta)$ is linear and sign-sensitive: a sign flip of any nonzero coordinate leaves $\|\theta\|_{1}$ unchanged but alters $L(\theta)$. Consequently, the techniques and results developed for norm estimation do not directly apply to linear functional estimation. In addition, the analysis in [8] does not accommodate non-Gaussian noise.
Theorem 1 generalizes the result in [9], which studied the homogeneous case where all $\eta_{j}=1$, to arbitrary loading vectors. In Section 5, we present explicit expressions for the minimax rate under several examples of loading vectors, including the homogeneous case.

### v2

@@ Abstract
We study estimation of the linear functional $\eta^{\top}\theta$ of a high-dimensional $s$-sparse mean vector $\theta$ when the loading vector $\eta$ is arbitrary and the noise is symmetric with exponentially decaying tails. Previous analyses for equal loadings treat coordinates as exchangeable and do not yield sharp rates when loadings vary. We give a sharp nonasymptotic characterization of the oracle minimax rate that makes explicit its dependence on $s$, $\eta$, and the noise tail parameter. To attain this rate, we construct an estimator that treats large and small loadings differently with a cutoff calibrated to $\eta$, and we prove a matching lower bound using a sparse prior whose inclusion probabilities and signal magnitudes depend on $\eta$. For unknown sparsity, we identify an $\eta$-dependent threshold for a Lepski type selection and show that the resulting estimator achieves the oracle minimax rate up to a logarithmic factor, and that it cannot be improved for a broad, verifiable class of loading vectors. In analytic examples, we demonstrate how heterogeneity in $\eta$ changes the minimax and adaptive rates. We also extend the theory to non-symmetric noise, hypothesis testing, and estimation with unknown noise variance.
@@ 1 Introduction
$L(\theta)=\eta^{\top}\theta=\sum_{j=1}^{d}\eta_{j}\theta_{j},$ (2)
based on the observations $y_{1},\ldots,y_{d}$, where $\eta=(\eta_{1},\ldots,\eta_{d})\in\mathbb{R}^{d}$ is a fixed loading vector. Because coordinates with zero loadings can be dropped, we assume that all $\eta_{j}$ ( $j=1,\ldots,d$) are nonzero and sorted such that $|\eta_{1}|\geq|\eta_{2}|\geq\cdots\geq|\eta_{d}|>0$.
@@
for some integer $s\in\{1,\ldots,d\}$. This sparsity assumption arises in many applications, including spectroscopy, astronomy, and interferometry [12]. In these applications, the observed signal is typically close to zero with a few rare spikes, often described as a nearly black object. Most theoretical work has focused on Gaussian or sub-Gaussian noise, although in practice other noise distributions may also be relevant.
In this paper, we denote by $\mathcal{P}_{j}$ the distribution of noise $\xi_{j}$ ($j=1,\ldots,d$) and by $\mathcal{P}_{\xi}=(\mathcal{P}_{1},\ldots,\mathcal{P}_{d})$ the joint noise distributions. We denote by $\mathbb{P}_{\theta,\mathcal{P}_{\xi}}$ the distribution of $(y_{1},\ldots,y_{d})$ when the signal is $\theta$ and the joint noise distribution is $\mathcal{P}_{\xi}$. We also write $\mathbb{E}_{\theta,\mathcal{P}_{\xi}}$ for the corresponding expectation. The classical Gaussian sequence model corresponds to the case where the noise variables $\xi_{i}$ are standard Gaussian, i.e., $\mathcal{P}_{1}=\cdots=\mathcal{P}_{d}=\mathcal{N}(0,1)$, in which case we write $\mathcal{P}_{\xi}=\mathcal{N}^{\otimes}$.
Beyond the Gaussian setting, we study a broader family of $\mathcal{P}_{\xi}$ that retains symmetry and fast-decaying tails. Specifically, we introduce the following class.
@@
Distributions in $\mathcal{G}_{\alpha,\tau}$ exhibit sub-Weibull tails and have recently attracted considerable attention [11, 23, 20]. The additional symmetry condition is important for our analysis. The class parameter $\alpha$ specifies the tail behavior, and the parameter $\tau$ specifies the spread of the tail. In particular, when $\alpha=2$, the class reduces to a sub-Gaussian class; when $\alpha=2$ and $\tau=2$, it includes the standard Gaussian distribution.
We denote by $\mathcal{G}_{\alpha,\tau}^{\otimes}$ the class of (independent) product distributions on $\mathbb{R}^{d}$ whose marginals all belong to $\mathcal{G}_{\alpha,\tau}$. In other words, $\mathcal{G}_{\alpha,\tau}^{\otimes}=\{\otimes_{j=1}^{d}P_{j}:P_{j}\in\mathcal{G}_{\alpha,\tau},~~j=1,\ldots,d\}$. We assume the noise distribution $\mathcal{P}_{\xi}$ lies in $\mathcal{G}_{\alpha,\tau}^{\otimes}$ with fixed $\alpha$ and $\tau$, but the marginals of $\mathcal{P}_{\xi}$ may differ and remain unknown.
@@
The minimax rate of estimation under the Gaussian sequence model has been extensively studied in the literature; see, for example, [15, 4, 5, 13, 14, 21]. The most closely related work to ours is [9], which established the nonasymptotic minimax rate for estimating the linear functional with a homogeneous loading vector (i.e., $\eta_{j}=1$ for all $j$) in the Gaussian sequence model. Specifically, for estimating the sum $\sum_{j=1}^{d}\theta_{j}$, they show that
@@
$\sigma^{2}s^{2}\log\!\left(1+\frac{d\log d}{s^{2}}\right).$
While the aforementioned theoretical works on the homogeneous loading case yield valuable insights into the roles of sparsity and dimensionality, they do not yield sharp minimax rates for estimating linear functionals with heterogeneous loadings (i.e., unequal $\eta_{j}$ values). Heterogeneous loadings commonly arise in practice, as illustrated in the following two examples.
@@
Let $\mathcal{H}$ be the class of square-integrable functions on $[0,1]$. Suppose $f\in\mathcal{H}$ is unknown and $Y$ is an observation from the white noise model
${\rm d\,}Y(t)=f(t)\,{\rm d\,}t+\sigma\,{\rm d\,}W(t)$
where $W(t)$ is a standard Brownian motion. It is often of interest to estimate the integral $\int fg$ for some test function $g\in\mathcal{H}$. To see the connection to our problem, let $\{\varphi_{i}(t)\}$ be any orthonormal basis for $\mathcal{H}$ and define
$y_{j}=\int_{0}^{1}\varphi_{j}\,{\rm d\,}Y,~~\theta_{j}=\int_{0}^{1}f\varphi_{j},~~\xi_{j}=\int_{0}^{1}\varphi_{j}\,{\rm d\,}W,~~\text{ and }\eta_{j}=\int_{0}^{1}\varphi_{j}g.$
It then follows that $\{\xi_{i}\}$ are independent standard normal r.v.s and $\int fg=\sum_{j}\eta_{j}\theta_{j}$. If it is believed that $\theta_{i}=0$ for $i>d$ and $\{\theta_{i}\}_{i=1}^{d}$ has at most $s$ nonzero components, this model coincides with (1) and the estimation reduces to estimating $L(\theta)$ in (2); see [17] for further discussions. In many problems, the test function $g$ is arbitrary, so the induced sequence $\{\eta_{j}\}$ may lack any assumed structural properties. ∎
@@
Given $n$ covariate vectors $x_{i}\in\mathbb{R}^{d}$, suppose we observe $z_{i}=x_{i}^{\top}\beta+\sigma\varepsilon_{i}$ where $\beta\in\mathbb{R}^{d}$ is the unknown parameter and $\varepsilon_{i}$ are i.i.d. samples from $\mathcal{N}(0,1)$. A common task is to predict the response for a new fixed covariate vector $x_{0}$ by estimating $x_{0}^{\top}\beta$. To see the connection to our problem, consider an orthogonal design, where the design matrix $\mathbf{X}$ can be written as $\mathbf{X}=\sum_{j\leq d}b_{j}u_{j}e_{j}^{\top}$, with singular values $b_{j}>0$, singular vectors $u_{j}\in\mathbb{R}^{n}$, and canonical basis vectors $e_{j}\in\mathbb{R}^{d}$. For any $j\leq d$, define the transformed variables and parameters as
$y_{j}=\sum_{i\leq n}u_{j,i}z_{i},~~\theta_{j}=b_{j}\beta_{j},~~\xi_{j}=\sum_{i\leq n}u_{j,i}\varepsilon_{i},\text{ and }~~\eta_{j}=b_{j}^{-1}x_{0,j}.$
It follows that $y_{j}=\theta_{j}+\sigma\xi_{j}$, $\xi_{j}$ are i.i.d. from $\mathcal{N}(0,1)$, and $x_{0}^{\top}\beta=\sum_{j=1}^{d}\eta_{j}\theta_{j}=L(\theta)$. The standard assumption that $\beta$ is $s$-sparse is now equivalent to $\theta$ being $s$-sparse. In this case, the model coincides with the sequence model in (1), and the prediction reduces to estimating $L(\theta)$, where the loading vector $\eta=(b_{j}^{-1}x_{0,j})$ is determined by an arbitrary $x_{0}$ and non-uniform singular values $b_{j}$. ∎
In both examples and many other applications, the loading vector is arbitrary and heterogeneous. This presents a critical gap in the existing minimax theories: the sequence model literature [4, 9, 10] assumes equal loadings, while the linear model literature [2, 6] considers special loading structures (specifically, loadings with bounded coordinate ratios or polynomial decay rates). Consequently, the optimal rates for general $\eta$ remain unknown. To our knowledge, the current paper is the first to address this problem.
Contributions. We study estimation of $L(\theta)=\eta^{\top}\theta$ for $s$-sparse $\theta$ with an arbitrary loading vector $\eta$ under any noise distribution $\mathcal{P}_{\xi}\in\mathcal{G}_{\alpha,\tau}^{\otimes}$. All results are nonasymptotic and explicit in $(s,d,\eta,\alpha,\sigma)$.
In the oracle setting where $s$ is known, we establish the minimax rate $\Phi_{o}(s;\eta)$, which is governed by an important quantity $\lambda_{o}$ that depends on $(s,\eta,\alpha)$ in Equation 6. Our theory not only recovers the homogeneous case, but also reveals how heterogeneity in $\eta$ changes the minimax rate. To obtain sharp upper and lower bounds, the derivation should take into account the structure of the loading vector. However, this aspect is completely missing in existing works on homogeneous loadings [4, 9, 10]. To fill this gap, we make the following technical contributions:
  • Upper bounds. The estimator in Equation 4 for homogeneous loadings is no longer optimal for heterogeneous loadings. To achieve minimax optimality with general $\eta$, it is necessary to treat large and small loadings differently. We use $\lambda_{o}$ as the cutoff value for $|\eta_{j}|$ and propose a hybrid estimator that handles $\theta_{j}$ with large $|\eta_{j}|$ by plug-in and those with small $|\eta_{j}|$ by thresholding. The cutoff and the threshold are precisely calibrated to the loading vector $\eta$ and the noise tail parameter $\alpha$; these choices are critical for achieving minimax optimality.
  • Lower bounds. The lower bound arguments in [9] are not sharp for heterogeneous loadings, as they rely on a uniform prior distribution that ignores the differences in $|\eta_{j}|$. Furthermore, their analysis is confined to Gaussian noise. To address these limitations, we develop new techniques that apply to general noise distributions in $\mathcal{G}_{\alpha,\tau}$ and construct a non-uniform sparse prior where the nonzero probabilities and the magnitudes of coordinates are specifically designed based on $\eta$ and $\alpha$. This refined construction yields sharp minimax lower bounds for general loading vectors and a broader family of noise distributions.
For unknown $s$, we investigate the price of adaptation. Unlike the Lepski type procedure for homogeneous loadings in [10], where the threshold for each candidate $s$ is independent of $\eta$ and is not optimal, we identify a new $\eta$-dependent threshold such that the resulting adaptive estimator attains a rate that matches the minimax rate up to a logarithmic factor in $s$. We further show that this rate is the unimprovable adaptive rate for a broad class of loading vectors satisfying a mild and verifiable condition (1).
@@
In addition, we study non-symmetric noise, showing that skewness can increase the minimax rate in certain examples. We also consider extensions to the setting with unknown noise variance $\sigma^{2}$ and to hypothesis testing for the linear functional.
@@
Organization. The rest of the paper is organized as follows. In Section 2, we study minimax estimation of the linear functional $L(\theta)$ under symmetric sub-Weibull noise when the sparsity level $s$ is known. Section 3 considers the more realistic case where $s$ is unknown and establishes the optimal adaptive rate for estimating $L(\theta)$. In Section 4, we extend the analysis to general (not necessarily symmetric) sub-Weibull noise with known sparsity. Section 5 provides analytically tractable examples of the derived rates. Extensions and future directions are discussed in Section 6 and Section 7, respectively. The last section contains the proofs of the lower bounds. Other proofs are provided in the supplementary material.
@@ 2 Minimax estimation with known sparsity
We first verify that $\beta$ is well defined. Indeed, the left-hand side of Equation 5 is continuous and strictly decreasing (see Lemma 11 in the supplementary material), diverges to $\infty$ as $\beta\to$-$\infty$, and converges to 0 as $\beta\to+\infty$. It follows that Equation 5 admits a unique solution $\beta$ and we have $\beta_{+}=\beta\vee 0\geq 0$. When $\alpha=2$, Equation 5 coincides with Equation (5) in [8], so our formulation extends their definition from Gaussian noise to symmetric sub-Weibull noise. The parameter $\lambda_{o}$ captures the dependence of the linear functional estimation problem on the three key components: the sparsity level $s$, the loading vector $\eta$, and the noise tail parameter $\alpha$. $\lambda_{o}$ plays a crucial role in characterizing the minimax rate, as it appears both in the expression of our lower bound and in the construction of our optimal estimator.
@@
and the estimator $\hat{L}_{s}$ is a minimax rate-optimal estimator.
@@
(1). [Lower bound] For $\alpha>0$, there exist constants $c,\tau_{\alpha}>0$ such that for any $\tau\geq\tau_{\alpha}$ and any integer $s\in[d]$, it holds that
@@
(2). [Upper bound] For the estimator $\hat{L}_{s}$ defined in (7) and $\alpha,\tau>0$, there exists a constant $C>0$ such that
@@
Theorem 1 may seem related to Theorem 1 of [8], who investigated the minimax separation distance in the heteroscedastic Gaussian sequence model for testing
$H_{0}:\;\theta=0\quad\text{against}\quad H_{1}:\;\|\theta\|_{t}\geq\varepsilon,$
where $\|\theta\|_{t}=(\sum_{j=1}^{d}|\theta_{j}|^{t})^{1/t}$ for $t\geq 1$. As discussed in their work, the theory of testing in this framework is closely connected to minimax estimation of the norm $\|\theta\|_{t}$. However, the case $t=1$ considered there fundamentally differs from our setting. The functional $\|\theta\|_{1}=\sum_{j=1}^{d}|\theta_{j}|$ is nonlinear and invariant to sign changes, whereas our target $L(\theta)$ is a linear and sign-sensitive functional: flipping the sign of any nonzero coordinate leaves $\|\theta\|_{1}$ unchanged but alters $L(\theta)$. This structural difference implies that the analytical tools and lower bound constructions used in norm estimation problems cannot be directly extended to our setting. In addition, the analysis in [8] does not accommodate non-Gaussian noise.
Theorem 1 extends the seminal result of [9], which focused on the homogeneous setting where all $\eta_{j}=1$, to the more general and practically relevant case of arbitrary loading vectors. This generalization reveals how heterogeneity in the loadings fundamentally alters the minimax separation rate and highlights the role of $\eta$ in governing the intrinsic difficulty of linear functional estimation. To illustrate these insights concretely, we provide in Section 5 explicit expressions for the minimax rate under several representative examples of loading structures, including the classical homogeneous case.